# Enhancing Adversarial Transferability via Component-Wise Augmentation Method

### v1

@@ 4.4 Evaluations on Defense Method
To comprehensively assess the robustness of our proposed method against diverse defense mechanisms, we conducted experiments under five defense settings: Adversarial Training (AT), High-level representation Guided Denoiser (HGD), Neural Representation Purifier (NRP), Randomized Smoothing (RS), and Diffusion Purification (Diffure), with the results for NRP included in the appendix 6. As summarized in Tables 3 through 6, our method, particularly CWA, consistently achieves the highest attack success rates across most of defense settings, outperforming BSR and other baseline methods. For instance, under Diffure, it outperforms BSR by 6.9% in mean ASR and exceeding weaker baselines by more than 20%. Even under RS, a certified defense known for its robustness, our method maintains competitive ASRs (e.g., 28.7% for ResNet-101 and 29.5% for Visformer), showcasing its resilience. These results validate the robustness and effectiveness of our approach across diverse defenses.
@@ 4.5 Ablation Study
To further evaluate the performance improvements introduced by CWA, we conduct ablation studies on five key hyperparameters and one critical step: number of blocks $n$, maximum scale factor $s_{\text{max}}$, maximum rotation angle $r$, number of rotated blocks $k$, number of transformed copies $N$, and the key step pre-interpolation. All experiments are conducted using ResNet-101 as the surrogate model.
On the number of blocks $n$.
The impact of block partitioning is illustrated in Figure 3 (a). When $n=1$, CWA applies global operations to the entire image, which fails to sufficiently disrupt the attention heatmaps. Increasing $n$ to $2$ significantly improves the attack success rate , as the $2\times 2$ block division introduces localized transformations, encouraging the model to focus on diverse regions of the object. When $n>2$, performance begins to degrade due to increased information loss from operating on a larger number of blocks.
Attack ResNet-18 ResNet-101 ResNeXt-50 DenseNet-121 ViT PiT Visformer Swin
@@
Attack ResNet-18 ResNet-101 ResNeXt-50 DenseNet-121 ViT PiT Visformer Swin
@@
However, in contrast to BSR, the performance does not drop sharply, as our selective rotation method stabilizes the results. Consequently, we set $n=2$.
Figure 3: Attack success rates (%) of various models on the adversarial examples generated by CWA with different numbers of blocks, maximum resize rates, maximum rotate angles, numbers of rotation blocks and numbers of transformed copies. The adversarial examples are crafted using the ResNet-101 model and tested on seven other models under the black-box setting.
@@
As demonstrated in Figure 3 (c), we observe that gradually increasing $r$ improves ASR, but fluctuations occur when the angle exceeds 25°. Considering the experimental error margins, we interpret this fluctuation as insignificant, and thus we select $r=26^{\circ}$ as the optimal value.
@@
Figure 4: Attack success rates (%) of various models on the adversarial examples generated by CWA with and without Pre-interpolation. The surrogate model is ResNet-101.
@@ 5 Conclusion
Based on the observation that inconsistent attention distributions across models are a major barrier to adversarial example transferability, we propose CWA, a method utilizing interpolation and selective rotation. CWA effectively generates diverse transformed images, enriching the attention distributions of individual models and reducing discrepancies in gradient loss calculations across different models. Experimental results demonstrate that our approach not only surpasses existing SOTA methods in ASR but also achieves significantly lower standard deviation, highlighting its stability and generalization capabilities. Furthermore, CWA offers a fresh perspective for input transformation-based adversarial attacks.
@@
6 Extended Evaluations on Defense Methods
In this section, we present the performance of our method on NRP, as shown in Tables 7 and 8. Specifically, we apply NRP to process adversarial images generated by different defense methods, and subsequently evaluate their Attack Success Rate (ASR) using eight models. The results demonstrate that our method, CWA, achieves superior performance in terms of both ASR and standard deviation (std) when applied to CNN-based and transformer-based surrogate models. Notably, while maintaining a reduction in standard deviation, our approach consistently improves the ASR by an average of 5% over the previous state-of-the-art methods.
Model Attack ResNet-18 ResNet-101 ResNeXt-50 DenseNet-121 ViT PiT Visformer Swin Mean Std. Dev.
ResNet-18 DIM 98.9* 39.5 40.7 64.5 19.5 25.8 30.4 34.7 44.3 25.8
@@ 6 Extended Evaluations on Defense Methods
CWA 83.1 60.9 64.7 98.6* 35.1 46.2 52.9 54.8 62.0 20.4
Table 7: Attack success rates (%) of adversarial examples generated using various attack methods across eight classification models under NRP. The surrogate models are cnn-based. * indicates the surrogate model.
Model Attack ResNet-18 ResNet-101 ResNeXt-50 DenseNet-121 ViT PiT Visformer Swin Mean Std. Dev.
ViT DIM 52.5 34.3 34.1 47.9 70.9* 34.4 35.4 39 43.6 13.1
@@
CWA 70.9 54.6 57.7 71.5 45.1 62.4 62.6 83.1* 63.5 11.7
Table 8: Attack success rates (%) of adversarial examples generated using various attack methods across eight classification models under NRP. The surrogate models are transformer-based. * indicates the surrogate model.
7 More Analysis on Interpolation
To better understand the role of interpolation in affecting model attention distributions, we conducted experiments using various interpolation methods (e.g., bilinear, bicubic, nearest neighbor, and area interpolation) under different scaling factors. As shown in Figure 5, scaling factors below 1.0 tend to disperse attention across irrelevant regions, reducing focus on critical object areas. Conversely, scaling factors significantly above 1.0 excessively concentrate attention on limited regions. Moderate scaling factors, typically ranging from 1.0 to 1.8, yield the most balanced attention distributions, effectively redistributing focus across diverse object regions. For example, as depicted in Figure 5(a), a scaling factor of 1.0 primarily directs attention to a person’s back, while a factor of 1.4 shifts focus towards the hips, and 1.8 further moves attention to the left shoulder.
@@ 7 More Analysis on Interpolation
Figure 5: Heatmaps of Different Interpolation Rates Using ResNet-18. (a) Area interpolation; (b) Bicubic interpolation; (c) Bilinear interpolation; (d) Nearest-neighbor interpolation.

### v2

@@ 4.4 Evaluations on Defense Method
To comprehensively assess the robustness of our proposed method against diverse defense mechanisms, we conducted experiments under four defense settings: Adversarial Training (AT), High-level representation Guided Denoiser (HGD), Randomized Smoothing (RS), and Diffusion Purification (Diffure). More extensive evaluations can be found in Appendix A. As summarized in Tables 3 through 6, our method, particularly CWT, consistently achieves the highest attack success rates across most of defense settings, outperforming BSR and other baseline methods. For instance, under Diffure, it outperforms BSR by 8.1% in mean ASR and exceeding weaker baselines by more than 17%. Even under RS, a certified defense known for its robustness, our method maintains competitive ASRs (e.g., 28.7% for RN-101 and 29.5% for ViF-S), showcasing its resilience. These results validate the robustness and effectiveness of our approach across diverse defenses.
@@ 4.5 Ablation Study
To further evaluate CWT, we conduct ablation studies on five key hyperparameters and one critical step: number of blocks $n$, maximum scale factor $s_{\text{max}}$, maximum rotation angle $r$, number of rotated blocks $k$, number of transformed copies $N$, and the key step pre-interpolation. All experiments are conducted using RN-101 as the surrogate model.
Attack RN-18 ($\uparrow$) RN-101 ($\uparrow$) RX-50 ($\uparrow$) DN-121 ($\uparrow$) ViT-B ($\uparrow$) PiT-B ($\uparrow$) ViF-S ($\uparrow$) Swin-T ($\uparrow$)
@@
Attack RN-18 ($\uparrow$) RN-101 ($\uparrow$) RX-50 ($\uparrow$) DN-121 ($\uparrow$) ViT-B ($\uparrow$) PiT-B ($\uparrow$) ViF-S ($\uparrow$) Swin-T ($\uparrow$)
@@
On the number of blocks $n$.
The impact of block partitioning is illustrated in Figure 4 (a). When $n=1$, CWT applies global operations to the entire image, which fails to sufficiently disrupt the attention heatmaps. Increasing $n$ to $2$ significantly improves the attack success rate , as the $2\times 2$ block division introduces localized transformations, encouraging the model to focus on diverse regions of the object.
Figure 4: Attack success rates (%) of various models on the adversarial examples generated by CWT with different numbers of blocks, maximum resize rates, maximum rotate angles, numbers of rotation blocks and numbers of transformed copies. The adversarial examples are crafted using the RN-101 model and tested on seven other models under the black-box setting.
When $n>2$, performance begins to degrade due to increased information loss from operating on a larger number of blocks. However, in contrast to BSR, the performance does not drop sharply, as our selective rotation method stabilizes the results. Consequently, we set $n=2$.
@@
As demonstrated in Figure 4 (c), we observe that gradually increasing $r$ improves ASR, but fluctuations occur when the angle exceeds 25°. Considering the experimental error margins, we interpret this fluctuation as insignificant, and thus we select $r=26^{\circ}$.
@@
Figure 5: Attack success rates (%) of various models on the adversarial examples generated by CWT with and without Pre-interpolation. The surrogate model is RN-101.
@@ 5 Conclusion
Based on the observation that inconsistent attention distributions across models are a major barrier to adversarial example transferability, we propose CWT, a method utilizing interpolation and selective rotation. CWT effectively generates diverse transformed images, enabling a single surrogate model to focus on different regions of the target object. Experimental results demonstrate that our approach not only surpasses existing SOTA methods in ASR but also achieves significantly lower standard deviation, highlighting its stability and generalization capabilities. Furthermore, CWT offers a novel perspective for input transformation-based adversarial attacks.
@@
Appendix A Extended Evaluations on Defense Methods
In this section, we evaluate the performance of our proposed method on five adversarially trained models, namely Inc-v3ens, Inc-v3ens3, Inc-v3ens4, IncRes-v2ens and Res50in  [38, 9], as well as the NRP [26] method. The results are summarized in Tables  A1 through  A3.
Specifically, in Table  A1, we analyze the attack effectiveness of adversarial samples generated using different attack methods on various surrogate models against adversarially trained models. It is worth noting that these models do not achieve 100% classification accuracy on clean images. However, the misclassification rates of the models are all below 5%. The results indicate that our method surpasses previous SOTA approaches by an average improvement of 5.3%.
For Tables A2 and  A3, we apply NRP to process adversarial images generated by different defense methods and ASR on eight models. The results demonstrate that our method, CWT, outperforms existing techniques in terms of both ASR and standard deviation for CNN-based and Transformer-based surrogate models. Notably, while maintaining a reduced standard deviation, our approach consistently improves the ASR by an average of more than 5% over prior SOTA methods.
Model Attack RN-18 ($\uparrow$) RN-101 ($\uparrow$) RX-50 ($\uparrow$) DN-121 ($\uparrow$) ViT-B ($\uparrow$) PiT-B ($\uparrow$) ViF-S ($\uparrow$) Swin-T ($\uparrow$) Mean ($\uparrow$) Std. Dev. ($\downarrow$)
Inc-v3ens DIM 64.2 50.7 46.2 68.0 47.8 48.8 61.4 52.4 54.9 7.8
SIM 60.1 45.2 41.2 67.2 49.0 44.1 54.8 28.6 48.8 11.2
Admix 71.6 58.9 50.6 77.9 51.4 43.8 59.5 31.1 55.6 14.0
MaskBlock 44.4 36.3 31.1 51.0 42.5 39.2 41.8 25.9 39.0 7.4
US-MM 68.7 60.1 52.8 74.4 53.5 47.2 64.8 30.8 56.5 12.9
BSR 88.9 79.1 71.5 89.8 72.7 75.0 81.4 75.8 79.3 6.5
CWT 92.7 84.6 79.0 92.6 74.0 82.1 86.1 84.0 84.4 5.9
Inc-v3ens3 DIM 60.2 48.3 43.7 65.7 44.7 47.1 58.4 49.8 52.2 7.6
SIM 57.1 43.7 36.4 63.8 46.7 41.2 51.7 26.1 45.8 11.1
Admix 67.9 56.4 49.7 75.1 49.5 42.1 59.2 29.3 53.6 13.5
MaskBlock 43.4 36.2 28.5 51.1 36.5 38.4 40.9 24.8 37.5 7.7
US-MM 65.0 60.1 51.2 72.3 50.1 45.2 61.9 29.8 54.5 12.5
BSR 85.7 77.7 70.5 87.1 71.0 71.6 78.5 73.6 77.0 6.1
CWT 91.4 82.5 77.5 90.8 72.5 79.4 85.5 81.6 82.7 6.0
Inc-v3ens4 DIM 61.5 45.2 39.1 65.2 43.9 46.0 54.8 48.7 50.6 8.5
SIM 52.7 39.8 34.4 62.8 44.5 38.2 48.6 25.8 43.4 10.7
Admix 64.3 53.2 45.7 72.9 47.5 40.0 54.6 28.0 50.8 13.1
MaskBlock 39.1 32.3 27.8 47.0 36.8 33.7 39.3 24.9 35.1 6.6
US-MM 61.6 56.8 46.8 71.0 49.7 43.5 58.5 27.5 51.9 12.4
BSR 85.1 73.6 66.8 86.4 69.2 70.5 77.2 72.9 75.2 6.8
CWT 90.9 82.0 75.4 90.8 72.1 79.1 81.7 79.9 81.5 6.2
IncRes-v2ens DIM 48.0 43.9 34.8 55.1 41.5 39.8 50.2 39.8 44.1 6.2
SIM 42.1 36.7 30.3 50.6 39.7 33.6 44.6 19.2 37.1 9.0
Admix 50.6 50.1 41.6 61.8 44.9 34.2 50.6 21.0 44.4 11.6
MaskBlock 30.1 28.4 23.3 37.7 33.6 29.2 30.0 17.5 28.7 5.7
US-MM 49.3 51.7 44.0 59.1 44.4 38.4 53.2 21.0 45.1 10.9
BSR 75.0 71.4 63.9 77.8 67.4 66.1 69.5 64.0 69.4 4.8
CWT 81.8 79.6 73.5 84.6 69.6 73.8 76.9 73.9 76.7 4.7
Res50in DIM 83.8 54.1 51.1 79.6 49.7 52.9 61.8 57.6 61.3 12.3
SIM 81.6 53.9 49.3 79.1 54.9 46.8 57.6 36.2 57.4 14.6
Admix 90.1 67.1 61.3 87.7 58.2 49.2 64.6 40.3 64.8 16.1
MaskBlock 72.4 47.1 39.8 70.5 48.3 44.7 47.6 33.3 50.5 13.0
US-MM 88.2 73.1 66.8 87.7 60.7 52.6 70.7 41.4 67.7 15.1
BSR 97.1 83.2 78.8 95.3 75.1 77.0 82.3 82.3 83.9 7.6
CWT 98.1 85.7 82.2 95.3 74.6 80.5 85.7 85.9 86.0 7.1
Table A1: Attack success rates (%) of adversarial examples generated using various attack methods across eight classification models under adversarially trained models.
Model Attack RN-18 ($\uparrow$) RN-101 ($\uparrow$) RX-50 ($\uparrow$) DN-121 ($\uparrow$) ViT-B ($\uparrow$) PiT-B ($\uparrow$) ViF-S ($\uparrow$) Swin-T ($\uparrow$) Mean ($\uparrow$) Std. Dev. ($\downarrow$)
RN-18 DIM 98.9* 39.5 40.7 64.5 19.5 25.8 30.4 34.7 44.3 25.8
@@ Appendix A Extended Evaluations on Defense Methods
CWT 83.1 60.9 64.7 98.6* 35.1 46.2 52.9 54.8 62.0 20.4
Table A2: Attack success rates (%) of adversarial examples generated using various attack methods across eight classification models under NRP. The surrogate models are CNN-based. * indicates the surrogate model.
Model Attack RN-18 ($\uparrow$) RN-101 ($\uparrow$) RX-50 ($\uparrow$) DN-121 ($\uparrow$) ViT-B ($\uparrow$) PiT-B ($\uparrow$) ViF-S ($\uparrow$) Swin-T ($\uparrow$) Mean ($\uparrow$) Std. Dev. ($\downarrow$)
ViT-B DIM 52.5 34.3 34.1 47.9 70.9* 34.4 35.4 39 43.6 13.1
@@
CWT 70.9 54.6 57.7 71.5 45.1 62.4 62.6 83.1* 63.5 11.7
Table A3: Attack success rates (%) of adversarial examples generated using various attack methods across eight classification models under NRP. The surrogate models are Transformer-based. * indicates the surrogate model.
Appendix B More Analysis on Interpolation
To better understand the role of interpolation in affecting model attention distributions, we conducted experiments using various interpolation methods (e.g., bilinear, bicubic, nearest neighbor, and area interpolation) under different scaling factors. As shown in Figure  A1, scaling factors below 1.0 tend to disperse attention across irrelevant regions, reducing focus on critical object areas. Conversely, scaling factors significantly above 1.0 excessively concentrate attention on limited regions. Moderate scaling factors, typically ranging from 1.0 to 1.8, yield the most balanced attention distributions, effectively redistributing focus across diverse object regions. For example, as depicted in Figure  A1 (a), a scaling factor of 1.0 primarily directs attention to a person’s back, while a factor of 1.4 shifts focus towards the hips, and 1.8 further moves attention to the left shoulder.
@@ Appendix B More Analysis on Interpolation
Figure A1: Heatmaps of Different Interpolation Rates Using ResNet-18. (a) Area interpolation; (b) Bicubic interpolation; (c) Bilinear interpolation; (d) Nearest-neighbor interpolation.
Appendix C Examples of Adversarial Images
In this section, we visualize the adversarial images generated by our method, CWT, on eight surrogate models, using heatmaps on ResNet-18, as shown in Figure A2. It can be observed that CWT effectively transfers the target attention to different regions of the image.
Figure A2: Heatmaps of the adversarial images generated by CWT on eight surrogate models. The heatmaps generated on ResNet-18.